if you invest $5,000 for a year and earn $400 invest,what is the annual interest rate ?( use the formula i=prt.)
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

I=PRT
400=5,000*R*1
400=5,000R
R=400/5,000
R=.08 OR 8% IS THE RATE OF INTEREST.

Answer by stanbon(75887)

if you invest $5,000 for a year and earn $400 invest,what is the annual interest rate ?( use the formula i=prt.)
----------
i=prt
400 = 5000*r*1
r = 400/5000
r = 0.08
r = 8%
============
Cheers,
Stan H.